OpenStudy (anonymous):

derivatives involving ln x: d/dx (lnx/lnx+1)

OpenStudy (freckles):

Apply quotient rule.

OpenStudy (anonymous):

\[((\ln x+1)(1/x)-lnx(1/x))/(lnx+1)^{2}\] \[(1/x)(\ln x+1-\ln x))/(\ln x+1)^{2}\] \[(1/x)(1)/(\ln x+1)^{2}\]